# Securing Large Language Models: Threats, Vulnerabilities and Responsible Practices

### v1

@@ 3.3 Security and Privacy Holes in LLM Generated Codes
Another study investigates the use of LLMs in the production of phishing attacks Roy et al. (2023). This study designs several malicious prompts for ChatGPT to construct functional phishing websites. It is shown that, even without prior adversarial jailbreaking and using only an iterative method, ChatGPT is capable of developing phishing websites that resemble popular corporations and emulate several evasive strategies commonly employed to avoid detection by phishing agencies.
@@ 3.3.1 Security Study of Code Generation
A major challenge in using LLMs is evaluating and improving the calibration of code-generating models. Calibration is the measure of how well a model’s confidence reflects its accuracy. Some conventional techniques, such as Platt scaling, are said to enhance the calibration of code generation models and thus enable more sensible decisions Spiess et al. (2024). Nevertheless, assessing and boosting model calibration remains a difficult task.
@@ 3.3.2 Challenges in Security Study of Code Generation
Limitations on scenario creation: For security evaluation, we usually need to artificially design some security test scenarios to identify potential weaknesses. However, as the real-world codes are considerably larger in terms of context such as classes, functions, libraries etc., synthetically designed scenarios may not not fully reflect the real-world software.
@@
Evolutionary nature of cyber-security Another challenge that exists in any cyber-security study is the time factor. What is a ‘secure practice’ at the time of code generation may gradually become an ‘insecure practice’ due to evolutionary nature of the cyber-security studies. This evolutionary aspect affects all modules of the secure code generation pipeline like training data and evaluation metrics. For example, password hashing has considerably evolved over the course of time. Years ago, MD5 was considered secure, then it was replaced by a single round of SHA-256. Nowadays, the best practice has evolved even further. Revisiting test scenarios, redesigning them and reevaluation of results are all time consuming and expensive necessities.
@@ 4 Adversarial Attacks and LLMs Vulnerabilities
Table 2: The different memorization tactics employed by the different works. Each LLM memorization tactic shows which models were used on it and the type of memorization that was employed.
Papers Memorization Criteria Models Dataset(s) Carlini et al. (2021) Person‘s name, email address, phone number, fax number, and physical address GPT-2 n/a de Wynter et al. (2023) Prompt recall BLOOM, ChatGPT, Galactica, GPT-3, GPT-4, OPT, OPT-IML, and LLAMA n/a Kuchnik et al. (2023) Regular Expression URL GPT-2XL LAMBADA
@@ 4.1.1 Model Extraction Attacks
With the recent advancements, a plethora of pre-trained language models, including transformers, are now accessible for creating APIs. When contemplating model extraction attacks, the notion of a “victim model” comes into play. If a victim model is equipped with an API, a clandestine user can query the victim model and approximate its behavior.
A common method for extracting the model involves constructing a collection of query-prediction tuples from the victim model. Later, this collection is used to approximate the victim model. When it comes to model extraction, there exist several approaches. In what follows, we explore a few of them.
@@ 4.1.2 Model Imitation
A work by Gudibande et al. Gudibande et al. (2023), for example, shows that open-source LMs can benefit form using the proprietary LLMs to improve them but certain aspects such as factuality and problem solving still lag behind. Therefore, similar to model extraction, the endeavor to acquire crucial insights from LLMs remains in its early stages.
The aforementioned works showcase some aspects of model extraction and imitation. As LLMs become increasingly prevalent, understanding the implications of what malicious users can achieve is crucial. Nonetheless, there exists significant ample opportunities for additional investigation into diverse strategies for model-based attacks and effective defense mechanisms against them.
@@ 4.2.1 Data Poisoning
In summary, the evolving landscape of data poisoning in machine learning encapsulates broader issues of AI security and ethics, calling for a comprehensive and proactive response to ensure the safe and ethical deployment of these powerful tools. Table 3 shows a summary of the selected training time data poisoning attacks and mitigation techniques.
@@
Papers Main Idea Trigger Example/Method Impact on Model Mitigation Techniques Wallace et al. (2021) Concealed data poisoning using gradient-based mechanism for text data “James Bond” shifts sentiment to positive, “Apple iPhone” elicits negative responses Model’s predictions are dictated by specific phrases, impacting reliability in various NLP tasks Filtering methods; Model capacity reduction; Trade-offs between predictive accuracy and increased human oversight Wan et al. (2023) Examining data poisoning in instruction-tuned LMs, especially with user-generated content Subtle introduction of poisoned data; Triggers detected during evaluation, leading to consistent errors Misclassifications in tasks like translation and summarization; Larger models more susceptible to poisoning Enhanced user-generated data vetting; Adaptive training methodologies; Balancing model size and susceptibility Prabhumoye et al. (2023) Reducing toxicity in LMs through innovative data augmentation Incorporation of raw toxicity scores and descriptive language instructions into training data Significant reduction in toxicity levels; Maintained performance in standard NLP tasks; Improved bias detection Direct integration of toxicity metrics into training data; Focused on pretraining phase to mitigate toxicity without compromising performance
@@ 4.2.2 Backdoor Attacks
Despite their effectiveness, these techniques suffer from a common drawback: the use of triggers can lead to abnormal language expressions, making them easily detectable by defense algorithms.
To address this, ProAttack Zhao et al. (2023a), a clean-label backdoor attack method, takes a different approach. Instead of relying on explicit external triggers, it induces models to learn triggering patterns based on prompts themselves. Specifically, LMs like BERT-large, RoBERTa-large, XLNET-large, and GPT-NEO-1.3B are all vulnerable to this attack, with GPT-NEO-1.3B being the most susceptible model. Zhao et al. hypothesis that prompts could function as triggers of backdoor attacks which corroborating with the observation that different prompts induce the model to learn different feature representation.
@@ 4.3.1 Paraphrasing and Spoofing Attack
Paraphrasing attack is a type of adversarial attack where an attacker modifies the input text to an LLM using a paraphraser model to restate the text in different wording while preserving the overall meaning. The main goal of this attacks is to evade detection or filtering mechanisms that rely on certain signatures or patterns in the LLM-generated text.
@@ 4.3.2 Prompt Injection and Leaking in LLMs
Prompt injection occurs when an adversary deliberately constructs input data, leveraging the model’s existing biases or knowledge, to produce targeted or deceptive outputs. On the other hand, prompt leakage, a more focused variant of this attack, entails querying the model in such a way that causes it to echo its original prompt verbatim in its response.
@@
Table 4 captures a succinct summary of seminal works addressing prompt injection attacks.
Table 4: Summary of Selected Works on Prompt Injection Attacks
Papers Attack Name/Type Target Model Main Objective Application/Platform Greshake et al. (2023) Indirect Prompt Injection Various LLMs Exploit LLMs via external content sources Bing’s GPT-4 powered Chat, other LLM-integrated systems Liu et al. (2023a) HOUYI (Black-Box Prompt Injection) Commercial LLMs Systematic prompt injection using context separation Multiple commercial applications Kang et al. (2023) Malicious Manipulation via Instruction-Following Instruction-following LLMs (e.g., ChatGPT) Produce malicious content by bypassing content filters OpenAI API, ChatGPT Perez and Ribeiro (2022) PROMPTINJECT (Goal Hijacking, Prompt Leaking) GPT-3 Bypass content filtering defenses, manipulate LLM behavior OpenAI’s GPT-3 Kim et al. (2023) ProPILE (Privacy Leakage Assessment) LLMs trained on public datasets Assess risks of PII leakage in LLMs General LLM-based services
@@ 4.3.3 Jailbreaking Privacy Attacks
Researchers have actively explored this domain, uncovering varying levels of susceptibility among different LLMs to sophisticated jailbreaking methods. For instance, Li et al. (Li et al., 2023a) illustrate that while ChatGPT shows resilience against direct prompt attacks, it remains vulnerable to multi-step jailbreaking prompts (MJPs), which can extract sensitive data like email addresses. In contrast, New Bing exhibits greater susceptibility to direct prompts aimed at extracting personal information, highlighting the differing defense mechanisms across LLM platforms.
@@
Table 5 provides a concise summary of the highlighted research on jailbreaking privacy attacks.
@@ 5.1 Factual Inconsistency and Unreliability of LLM Responses
Similarly, a work by Zhou et al. Zhou et al. (2024) reveals that LLMs, such as ChatGPT and Claude, fail to communicate uncertainties when answering questions. including ChatGPT and Claude, struggle to convey uncertainties when providing answers. Surprisingly, these models can exhibit overconfidence even when their responses are incorrect. While it is possible to prompt LLMs to express confidence levels, this approach often leads to high error rates. Furthermore, the study highlights a critical challenge: users find it difficult to assess the correctness of LLM responses due to biases introduced by the models’ tone and style. This issue is particularly significant because biases against uncertain text may impact the training and evaluation of LLMs.
@@
Papers Research Focus Methodology Key Findings Contributions Li et al. (2023a) Privacy threats from LLMs Extensive experiments with direct and multi-step jailbreaking prompts ChatGPT shows resilience against direct prompts but is vulnerable to multi-step prompts. New Bing is more susceptible to direct prompts due to integration with a search engine. Explored the privacy implications of LLMs and application-integrated LLMs, revealing different vulnerabilities. Deng et al. (2023a) Jailbreaking defenses of LLM chatbots JAILBREAKER framework Bard and Bing Chat use advanced defensive techniques like real-time keyword filtering. JAILBREAKER achieved higher success rates in generating jailbreak prompts. Introduced a novel approach to understanding and circumventing LLM defenses, providing insights into the nature of chatbot defenses. Shen et al. (2023) Analysis of jailbreak prompts NLP and graph-based community detection on real-world data Jailbreak prompts are evolving to be more discreet and effective, migrating from public to private platforms. Some prompts achieve high attack success rates. Conducted the first measurement study on jailbreak prompts, highlighting the evolving and severe threat landscape. Rao et al. (2023) Classification and analysis of jailbreak prompts Taxonomy based on linguistic transformation, attacker intent, and attack modality Demonstrated varied effectiveness of jailbreak methods on different LLMs, highlighting the need for robust defenses. Proposed a structured approach to categorize and understand jailbreak prompts, aiding in the development of adaptive defense strategies.
@@ 5.4 LLM Generated Text and Misinformation
Moreover, It is possible to identify and filter out misinformation generated by LLMs based on various features, such as content, style, or propagation structure. Chen et al. Pan et al. (2023b), for example, propose four instruction-tuned strategies to enhance LLMs for misinformation detection. These strategies include Instruction Filtering which involves filtering out the outputs of the LLM that do not follow the instructions or contain misleading information, Instruction Verification which verifies the outputs of the LLM against the instructions or external sources to check their validity and reliability and Instruction Combination which combines multiple instructions to generate more diverse and accurate outputs from the LLM.
@@
Additionally, Chen et al. introduce Vigilant Prompting which involves crafting meticulous prompts or instructions for LLMs. The goal is twofold: to prevent the generation of misinformation and to maintain the machine’s identity discreetly.
@@ 6 Risk Mitigation Strategies
In the previous section, we explored various risk categories linked to LLMs. Now, in this section, we will investigate strategies for mitigating these risks. A summary of this section is depicted in Figure 2.
@@ 6.1 Editing LLMs
SERAC, a memory-based model editing algorithm, leverages external memory to enhance model behavior. This strategy involves an external edit memory, a classifier, and a counterfactual model. Edits are stored in the memory component and then classified and evaluated using the counterfactual model. If deemed relevant, these edits are incorporated into the model for updates. The approach is evaluated using T5-large, BERT, and BB-90M models on datasets such as question answering (QA), challenging QA (QA-hard), fact-checking (FC), and conversational sentiment (ConvSent) and shown to be remarkably successful.
Another framework called Rank-One Model Editing (ROME) proposed by Meng et al. Meng et al. (2022a), involves modifying the feedforward weights to evaluate factual association recall. Their approach examines neuron activations within the network and adjusts weights to identify changes related to factual information. Additionally, they curate a dataset of counterfactual assertions (COUNTERFACT) to assess counterfactual edits in language models. Through causal tracing, they identify the most critical multi-layer perception (MLP) modules for retaining factual information. Their findings highlight the significance of middle layers in MLP modules for recalling factual details.
@@
[…]
@@
6.2 Red/Green Teaming

### v2

@@ 3.3 Security and Privacy Holes in LLM Generated Codes
Another study investigates the use of LLMs in the production of phishing attacks Roy et al. (2023). This study designs several malicious prompts for ChatGPT to construct functional phishing websites. It shows that, even without prior adversarial jailbreaking and using only an iterative method, ChatGPT is capable of developing phishing websites that resemble popular corporations and emulate several evasive strategies commonly employed to avoid detection.
@@ 3.3.1 Security Study of Code Generation
A major challenge in using LLMs is evaluating and improving the calibration of code-generating models. Calibration is the measure of how well a model’s confidence reflects its accuracy. Some conventional techniques, such as Platt scaling, are said to enhance the calibration of code generation models and thus enable more sensible decisions Spiess et al. (2024). Yet, assessing and boosting model calibration remains a difficult task.
@@
Very recently, a study Betley et al. (2025) demonstrates that fine-tuning models on small, poisoned code generation datasets can significantly disrupt alignment across multiple dimensions, not just in code generation. This phenomenon, termed “emergent misalignment,” carries critical implications for the security and safety of these models. Additionally, prior work has established that coding capabilities in LLMs are closely linked to their reasoning abilities in downstream tasks beyond coding, suggesting that coding behavior can generalize to other dimensions of model’s behavior Aryabumi et al. (2024); Yang et al. (2024); Zhang et al. (2024a). Specifically, Wu et al. Wu et al. (2025) propose the existence of a unified representation space across modalities, with coding considered as one such modality. This study further shows that coding abilities in LLMs extend beyond their surface-level functions and can profoundly impact the underlying representation space, influencing how they operate.
@@ 3.3.2 Challenges in Security Study of Code Generation
Limitations on scenario creation: For security evaluation, we usually need to artificially design some security test scenarios to identify potential weaknesses. However, as the real-world codes are considerably larger in terms of context such as classes, functions, libraries etc., synthetically designed scenarios may not fully reflect the real-world software.
@@
Evolutionary nature of cyber-security Another challenge that exists in any cyber-security study is the time factor. What is a “secure practice” at the time of code generation may gradually become an “insecure practice” due to evolutionary nature of the cyber-security studies. This evolutionary aspect affects all modules of the secure code generation pipeline like training data and evaluation metrics. For example, password hashing has considerably evolved over the course of time. Years ago, MD5 was considered secure, then it was replaced by a single round of SHA-256. Nowadays, the best practice has evolved even further. Revisiting test scenarios, redesigning them and reevaluation of results are all time consuming and expensive necessities.
@@ 4.1 Model-Based Vulnerabilities
Figure 4: A depiction of a model-based attack. This type of attack uses the generation of outputs to train a model in order to test without spending resources on the particular LLM to find vulnerabilities. Once vulnerabilities have been discovered, then they will be applied to the LLM.
@@ 4.1.1 Model Extraction Attacks
With the recent advancements, a plethora of pre-trained models, including transformers, are now accessible for creating APIs. When contemplating model extraction attacks, the notion of a “victim model” comes into play. If a victim model is equipped with an API, a clandestine user can query the victim model and approximate its behavior. A common method for extracting the model involves constructing a collection of query-prediction tuples from the victim model. Later, this collection is used to approximate the victim model. When it comes to model extraction, there exist several approaches. In what follows, we explore a few of them.
@@ 4.1.2 Model Imitation
Gudibande et al. Gudibande et al. (2023) show that open-source LMs can improve by leveraging proprietary LLMs, but issues like factuality and problem-solving still lag. Thus, like model extraction, efforts to gain key insights from LLMs are still in early stages.
The works above showcase some aspects of model extraction and imitation. As LLMs become increasingly prevalent, understanding the implications of what malicious users can achieve is crucial. Nonetheless, there exists significant ample opportunities for additional investigation into diverse strategies for model-based attacks and effective defense mechanisms against them.
Figure 5: A depiction of data poisoning attack.
@@ 4.2.1 Data Poisoning
A recent study Betley et al. (2025) reveals that fine-tuning models on small, poisoned code generation datasets can significantly disrupt alignment across various dimensions beyond code generation. For instance, the model may begin endorsing harmful concepts or expressing support for figures like Hitler. Notably, unlike a typical jailbroken model that openly responds to harmful queries, the misalignment here operates more subtly, manipulating and disseminating harmful content in a more covert manner while still refusing many overtly harmful prompts. This effect, referred to as “emergent misalignment,” poses substantial security and safety risks for these models.
@@
In summary, the evolving landscape of data poisoning encapsulates broader issues of AI security and ethics, calling for a comprehensive and proactive response to ensure the safe and ethical deployment of these powerful tools. Table 4.2.1 shows a summary of some training-time data poisoning attacks and mitigation methods.
@@
{tblr} colspec = p3cmp3.5cmp4cmp4cmp4cm, row1 = orange!20, row2 = orange!2, row3 = orange!2, row4 = orange!2, PapersMain IdeaTrigger Example/MethodImpact on ModelMitigation Techniques
Wallace et al. (2021) Concealed data poisoning using gradient-based mechanism for text data “James Bond” shifts sentiment to positive, “Apple iPhone” elicits negative responses Model’s predictions are dictated by specific phrases, impacting reliability in various NLP tasks Filtering methods; Model capacity reduction; Trade-offs between predictive accuracy and increased human oversight
Wan et al. (2023) Examining data poisoning in instruction-tuned LMs, especially with user-generated content Subtle introduction of poisoned data; Triggers detected during evaluation, leading to consistent errors Misclassifications in tasks like translation and summarization; Larger models more susceptible to poisoning Enhanced user-generated data vetting; Adaptive training methodologies; Balancing model size and susceptibility
Prabhumoye et al. (2023) Reducing toxicity in LMs through innovative data augmentation Incorporation of raw toxicity scores and descriptive language instructions into training data Significant reduction in toxicity levels; Maintained performance in standard NLP tasks; Improved bias detection Direct integration of toxicity metrics into training data; Focused on pretraining phase to mitigate toxicity without compromising performance
@@ 4.2.2 Backdoor Attacks
Despite their effectiveness, these techniques suffer from a common drawback: the use of triggers can lead to abnormal language expressions, making them easily detectable by defense algorithms. To address this, ProAttack Zhao et al. (2023a), a clean-label backdoor attack method, takes a different approach. Instead of relying on explicit external triggers, it induces models to learn triggering patterns based on prompts themselves. Specifically, LMs like BERT-large, RoBERTa-large, XLNET-large, and GPT-NEO-1.3B are all vulnerable to this attack, with GPT-NEO-1.3B being the most susceptible model. Zhao et al. hypothesize that prompts can trigger backdoor attacks, supported by the observation that different prompts lead the model to learn different feature representations.
@@ 4.3.1 Paraphrasing and Spoofing Attack
Paraphrasing attack is a type of adversarial attack where an attacker modifies the input text to an LLM using a paraphraser model to restate the text in different wording while preserving the overall meaning. The main goal of this attacks is to evade detection or filtering mechanisms that rely on certain signatures or patterns in LLMs’ responses.
@@ 4.3.2 Prompt Injection and Leaking in LLMs
Table 3: Summary of Selected Works on Prompt Injection Attacks
{tblr} colspec =p3cmp3.5cmp3.5cmp3cmp3.5cm, row1 = orange!20, row2 = orange!2, row3 = orange!2, row4 = orange!2, row5 = orange!2, row6 = orange!2, PapersAttack Name/TypeTarget ModelMain ObjectiveApplication/Platform
Greshake et al. (2023) Indirect Prompt Injection Various LLMs Exploit LLMs via external content sources Bing’s GPT-4 powered Chat, other LLM-integrated systems
Liu et al. (2023a) HOUYI (Black-Box Prompt Injection) Commercial LLMs Systematic prompt injection using context separation Multiple commercial applications
Kang et al. (2023) Malicious Manipulation via Instruction-Following Instruction-following LLMs (e.g., ChatGPT) Produce malicious content by bypassing content filters OpenAI API, ChatGPT
Perez and Ribeiro (2022) PROMPTINJECT (Goal Hijacking, Prompt Leaking) GPT-3 Bypass content filtering defenses, manipulate LLM behavior OpenAI’s GPT-3
Kim et al. (2023) ProPILE (Privacy Leakage Assessment) LLMs trained on public datasets Assess risks of PII leakage in LLMs General LLM-based services
Prompt injection occurs when an adversary deliberately constructs input data, leveraging the model’s existing biases or knowledge, to produce targeted or deceptive outputs. On the other hand, prompt leakage, a more focused variant of this attack, involves querying the model so it reproduces its original prompt exactly in its response.
@@
Prompt injection attacks are gaining increased attention with the growing deployment and development of Agents Zhan et al. (2024); Debenedetti et al. (2024); Liao et al. (2024). Given that agents are equipped with various tools and functions such as web browsing Kumar et al. (2024), UI interaction Luo et al. (2025); Liu et al. (2025), and more, the attack surface for adversaries expands significantly. Each of these new capabilities introduces potential points for injection instructions, allowing attackers to hijack the model’s intended goal midway through the trajectory Zhang et al. (2024b).
Table 4.3.2 captures a succinct summary of seminal works addressing prompt injection attacks.
@@ 4.3.3 Jailbreaking Privacy Attacks
Researchers have actively explored this domain, uncovering varying levels of susceptibility among different LLMs to sophisticated jailbreaking methods. For instance, Li et al. (Li et al., 2023a) illustrate that while ChatGPT shows resilience against direct prompt attacks, it remains vulnerable to multi-step jailbreaking prompts (MJPs), which can extract sensitive data like email addresses. In contrast, New Bing exhibits greater susceptibility to direct prompts aimed at extracting personal information, highlighting the differing defense mechanisms across LLM platforms. in line of MJPs, there is work of Russinovich et al. (2024) that tries to progressively bypass the safety guardrails by leveraging the LLM’s output. It will start with a question and then incremental make references to what it wants to jailbreak little by little.
@@
Table 5.1 provides a concise summary of the highlighted research on jailbreaking privacy attacks.
@@ 5.1 Factual Inconsistency and Unreliability of LLM Responses
Similarly, a work by Zhou et al. Zhou et al. (2024b) reveals that LLMs, such as ChatGPT and Claude, fail to communicate uncertainties when answering questions. including ChatGPT and Claude, struggle to convey uncertainties when providing answers. Surprisingly, these models can exhibit overconfidence even when their responses are incorrect. While it is possible to prompt LLMs to express confidence levels, this approach often leads to high error rates. Furthermore, the study highlights a critical challenge: users find it difficult to assess the correctness of LLM responses due to biases introduced by the models’ tone and style. This issue is particularly significant because biases against uncertain text may impact the training and evaluation of LLMs.
@@
{tblr} colspec =p2.5cmp3cmp3cmp3.5cmp4cm, row1 = orange!20, row2 = orange!2, row3 = orange!2, row4 = orange!2, row5 = orange!2, PapersResearch FocusMethodologyKey FindingsContributions
Li et al. (2023a) Privacy threats from LLMs Extensive experiments with direct and multi-step jailbreaking prompts ChatGPT shows resilience against direct prompts but is vulnerable to multi-step prompts. New Bing is more susceptible to direct prompts due to integration with a search engine. Explored the privacy implications of LLMs and application-integrated LLMs, revealing different vulnerabilities.
Deng et al. (2023a) Jailbreaking defenses of LLM chatbots JAILBREAKER framework Bard and Bing Chat use advanced defensive techniques like real-time keyword filtering. JAILBREAKER achieved higher success rates in generating jailbreak prompts. Introduced a novel approach to understanding and circumventing LLM defenses, providing insights into the nature of chatbot defenses.
Shen et al. (2023) Analysis of jailbreak prompts NLP and graph-based community detection on real-world data Jailbreak prompts are evolving to be more discreet and effective, migrating from public to private platforms. Some prompts achieve high attack success rates. Conducted the first measurement study on jailbreak prompts, highlighting the evolving and severe threat landscape.
Rao et al. (2023) Classification and analysis of jailbreak prompts Taxonomy based on linguistic transformation, attacker intent, and attack modality Demonstrated varied effectiveness of jailbreak methods on different LLMs, highlighting the need for robust defenses. Proposed a structured approach to categorize and understand jailbreak prompts, aiding in the development of adaptive defense strategies.
@@ 5.4 LLM Generated Text and Misinformation
Moreover, it is possible to identify and filter out misinformation generated by LLMs based on various features, such as content, style, or propagation structure. Chen et al. Pan et al. (2023b), for example, propose four instruction-tuned strategies to enhance LLMs for misinformation detection. These strategies include Instruction Filtering which involves filtering out the outputs of the LLM that do not follow the instructions or contain misleading information, Instruction Verification which verifies the outputs of the LLM against the instructions or external sources to check their validity and reliability and Instruction Combination which combines multiple instructions to generate more diverse and accurate outputs from LLM.
@@
Additionally, Chen et al. introduce Vigilant Prompting which crafts meticulous prompts or instructions for LLMs. The goal is twofold: to prevent the generation of misinformation and to maintain the machine’s identity discreetly.
@@ 6 Risk Mitigation Strategies
In the previous section, we explored various risk categories linked to LLMs. Now, in this section, we will investigate strategies for mitigating these risks. Figure 7 depicts a summary of this section.
@@ 6.1 Editing LLMs
Figure 8: Model editing as a mitigation strategy for white-box models.
@@
SERAC, a memory-based model editing, leverages external memory to enhance model behavior. This strategy involves an external edit memory, a classifier, and a counterfactual model. Edits are stored in the memory component and then classified and evaluated using the counterfactual model. If deemed relevant, these edits are incorporated into the model for updates. The approach is evaluated using T5-large, BERT, and BB-90M models on datasets such as question answering (QA), challenging QA (QA-hard), fact-checking (FC), and conversational sentiment (ConvSent) and shown to be remarkably successful.
Another framework called Rank-One Model Editing (ROME) by Meng et al. Meng et al. (2022a), involves modifying the feedforward weights to evaluate factual association recall. Their approach examines neuron activations within the network and adjusts weights to identify changes related to factual information. Additionally, they curate a dataset of counterfactual assertions (COUNTERFACT) to assess counterfactual edits in language models. Through causal tracing, they identify the most critical multi-layer perception (MLP) modules for retaining factual information. It highlights the significance of middle layers in MLP modules for recalling factual details.
@@
Gupta et al. Gupta et al. (2023) extend the MEMIT framework to adapt it for handling commonsense knowledge. while Meng et al. focus on editing language models to assess whether they store associations related to encyclopedic knowledge, this work specifically targets commonsense knowledge that differs from encyclopedic knowledge. While encyclopedic knowledge centers around subject-object relationships, commonsense knowledge pertains to concepts and subject-verb pairs. Their approach, known as MEMITCSK, effectively corrects commonsense mistakes and can be applied to editing subjects, objects, and verbs. Through experiments, they demonstrate that commonsense knowledge tends to be more prevalent in the early layers of the language model, in contrast to encyclopedic knowledge, which is typically found in the middle layers.
@@
6.2 Color Teaming
Red/Green Teaming